# QuickVideo: Real-Time Long Video Understanding with System Algorithm Co-Design

### v1

@@ 3 Method
In this section, we first introduce QuickVideo, mainly consisting of three components:
@@ 3.1 QuickCodec: Long Video Decoding for VideoLLMs
Given a bit stream $\mathcal{S}$ for a video $\mathcal{V}=(f_{1},f_{2},\dots,f_{m})$ with frame height $h$, frame width $w$ and a desired degree of concurrency $c$, our goal is to compute ${\bm{F}}$ such that $\forall j\in\{0,1,\dots,|\mathcal{I}|-1\}\;{\bm{F}}_{j}=f_{\mathcal{I}[j]}$ for $\mathcal{I}\subseteq\{1,2,\dots,m-1\}$. That is, ${\bm{F}}$ is a packed tensor containing all the frames selected in $\mathcal{I}$. We assume that $m$ is known from container metadata or an estimate using $pts_{max}$ and $pts_{min}$. The efficiency of our algorithm relies on two observations:
(1) It is faster to use $c$ cores to decode $c$ short videos than use $c$ cores to sequentially decode $1$ long video. Video decoding for human playback focuses on the second case, as humans watch earlier frames while later frames decode. However, due to interframe dependencies, sequential video decoding is difficult to parallelize [38]. Unlike in human playback, VideoLLMs require the entire video input to be loaded upfront. Therefore, we can decompose the loading of a long video $\mathcal{V}$ into loading $c$ short videos that span $\mathcal{V}$. However, we cannot begin video decoding from an arbitrary frame, only a keyframe. Keyframe Intervals (Appendix A) is a subroutine that parses the metadata of $\mathcal{S}$ and computes $c$ approximately equal length intervals, starting and ending on keyframes, that span $\mathcal{V}$. We parallelize over these intervals in Algorithm 2.
(2) Ideally, VideoLLMs sample frames at a short, regular interval, usually 1-2 FPS [4]. This is less than the gap between keyframes in standard codecs. Therefore, seek-based decoding must decode from all keyframes regardless, and many seeks are redundant. Our algorithm only requires $1$ seek operation per core, instead of seeks proportional to the number of frames sampled.
@@
Algorithm 2 describes the core of our video decoding algorithm. The algorithm begins by using metadata to calculate $c$ keyframe-aligned intervals span the video (line 1). Lines 2-5 initialize a block of shared memory ${\bm{F}}$ and compute a dictionary $M$ that maps indices of selected frames to unique memory offsets in ${\bm{F}}$. We then decode the long video in $c$ parallel intervals (lines 6-19). Video decoding begins by seeking the the start of the interval $pts_{start}$, which is guaranteed to be a keyframe (line 7). We enqueue packet to decode (lines 17-18) until the video decoder yields frames to process (line 9). If the timestamp of the dequeued frame is greater than or equal to the endpoint of the interval $pts_{end}$, parallel processing ends (lines 11-12 and 19). As the intervals in $\mathcal{J}$ span $\mathcal{S}$, $pts_{min}$ and $pts_{max}$ are given by the least and greatest values in $\mathcal{J}$, respectively. Therefore, we use can Equation 1 to calculate the index $i$ of $f$ (line 13). Lastly, we save $f$ to ${\bm{F}}$ if $f$ is a selected frame (lines 14-16). As decoding from a keyframe yields all frames with greater $pts$ and $\mathcal{D}$ outputs frames in $pts$ order, when the parallelized loop exits (line 19), all selected frames with $pts$ in the interval $[pts_{start},pts_{end})$ will have been output by $\mathcal{D}$ and saved to ${\bm{F}}$. Therefore, as $\mathcal{J}$ spans $\mathcal{S}$, when the algorithm returns, ${\bm{F}}$ will contain all selected frames.
3.2 QuickPrefill: Efficient Group-based Pre-filling for VideoLLM
@@
After decoding the video bit stream into packed tensors, they are then fed into the VideoLLM for inference. However, LLM generation with a long context is a well-known challenge due to its intensive memory usage and costly computation. To mitigate this issue, we introduce QuickPrefill, a grouped prefilling and KV cache pruning method that both accelerates the speed and reduces the memory requirement significantly.
Group-based Prefill.
@@
Let $V={\mathbf{v}_{1},\mathbf{v}2,\ldots,\mathbf{v}{|\mathbf{X}^{v}|}}$ denote the sequence of video tokens, where $|\mathbf{X}^{v}|$ is the total number of tokens and each token $\mathbf{v}_{i}\in\mathbb{R}^{d}$ is a $d$-dimensional vector. To reduce memory overhead during prefilling, we adopt a group-based strategy by partitioning the video token sequence into $G$ disjoint groups: $\mathbf{V}={\mathbf{V}_{1},\ldots,\mathbf{V}_{G}}$, where each group $\mathbf{V}_{g}$ contains approximately $N_{g}=\frac{|\mathbf{X}^{v}|}{G}$ tokens. Instead of processing the entire sequence at once, we sequentially prefill each group and store its corresponding key-value (KV) cache as $\mathbf{K}_{g}^{(l)},\mathbf{V}_{g}^{(l)}$ for each transformer layer $l$. This strategy significantly reduces the peak activation memory usage by a factor of $G$ and remains effective even when used in conjunction with efficient attention mechanisms such as FlashAttention. Empirically, it enables hour-long video understanding while keeping GPU memory usage within practical limits (e.g., reducing more than 100 GB of memory overhead; see subsection D.2).
Group-based KV Cache Pruning.
@@
While group-based prefill can effectively reduce the peak activation memory, the largest memory bottleneck, KV cache memory is still not resolved. Therefore, when processing each group of video tokens, instead of saving the all KV cache vectors, we will prune unimportant KV cache to maintain an $\rho\in(0,1]$ retention ratio and thus reduce the KV cache memory usage by a factor of $\frac{1}{\rho}$. The pruning decision is based on an importance score function $\mathbf{s}$ to produce an ordered list where we select the top-k KV cache until the retention ratio is reached:
@@
$\tilde{\mathbf{K}}_{g}^{(l)}=\mathbf{K}_{g}^{(l)}[I_{g}^{(l)}],\text{ }\tilde{\mathbf{V}}_{g}^{(l)}=\mathbf{V}_{g}^{(l)}[I_{g}^{(l)}],\text{ where }I_{g}^{(l)}=\operatorname{TopK}\left(\mathbf{s}(\mathbf{K}_{g}^{(l)},\mathbf{V}_{g}^{(l)}),\;k=\rho\cdot N_{g}\right)$ (2)
@@
where $\operatorname{TopK}$ defines a function that returns of the indices the top-k largest inputs. There are multiple heuristic importance score function $\mathbf{s}$ raised in previous works [11, 16, 45], in this paper, we mainly use the following 3 importance function: 1) Key Norms (small): $\mathbf{s}=-L_{2}(\mathbf{K}_{g}^{(l)})$; 2) Value Norms: $\mathbf{s}=L_{2}(\mathbf{V}_{g}^{(l)})$; 3) Attention Scores: $\mathbf{s}=\text{matmul}(\mathbf{K}_{g}^{(l)},\mathbf{Q}^{(l)})$. Here $L_{2}$ denotes the L2-norm function and $\mathbf{Q}^{(l)}\in\mathbb{R}^{|\mathbf{X}^{t}|\times(n_{h}\times d_{h})}$ denotes query vector of the text tokens in layer $l$. For our QuickPrefill, we use Key Norms (small) as the default importance function, due to its good performance.
@@ 3.3 Overlapping QuickCodec and QuickPrefill
The preceding sections introduced two complementary components: QuickCodec for CPU-based video decoding and QuickPrefill for GPU-based group-wise prefilling. However, sequential execution of these components leads to suboptimal resource utilization; the GPU remains idle during video decoding while the CPU is underutilized during prefill. To address this inefficiency, we propose an overlapped execution scheme that enables concurrent processing across CPU and GPU resources.
Overlapping QuickCodec and QuickPrefill requires a small adaptation to how frames are loaded. Instead of using $c$ cores to load $c$ intervals, we divide $\mathcal{V}$ into $s$ intervals, where $s\gg c$ using Keyframe Intervals($\mathcal{V},s$). We then load the frames from the $s$ intervals using $c$ cores, where the intervals corresponding to earlier blocks of video are loaded first. This approach allows us to use QuickCodec’s fast video decoding, while ensuring that early frames in $\mathcal{V}$ are prioritized, allowing QuickPrefill to begin prefilling groups on GPU. Once the video frames required for the first group are loaded, QuickPrefill immediately begins prefilling. QuickCodec continues to load frames using the CPU in the background. After QuickPrefill finishes prefill for a group, it saves the resulting KV cache vectors and checks if QuickCodec has loaded the frames required to process the next block. If the frames are available, QuickPrefill can begin processing the next group immediately. This establishes a producer-consumer pipeline between CPU video decoding and GPU prefill, where the GPU is only idle if the frames for its next block are not loaded.
The performance benefit of this overlapping execution can be formally analyzed. Let $t_{dec}$ and $t_{prefill}$ denote the total time for decoding and prefilling all video groups, respectively. In the sequential approach, the total execution time is $t_{dec}+t_{prefill}$. However, with our overlapped strategy, the execution time reduces to:
$t_{total}=\max\{t_{dec}+t^{g}_{prefill},t_{prefill}+t^{g}_{dec}\}+\Delta$ (3)
Where $t^{g}_{dec}$ and $t^{g}_{prefill}$ represent the time to decode frames for the first group and prefill the last group, respectively. $\Delta$ is a small amount latency added by QuickCodec parsing the video metadata. As an individual group corresponds to a small number of frames and video tokens, this approach achieves near-optimal overlap between CPU and GPU resources. This results in substantial speedup for hour-scale video processing. We note that some VideoLLMs have additional input preprocessing steps, such as calculating position embeddings or normalization, which we exclude from this analysis [4].
@@ 4 Experiments
We evaluate QuickVideo’s performance on practical long video understanding tasks. In section 4.1, we benchmark QuickCodec against existing frameworks. We also examine the limitations of QuickCodec, identifying use-cases where seek-based frameworks (Algorithm 1) have stronger performance. Next, in section 4.2 we evaluate the performance of QuickPrefill on four long video understanding benchmarks and analyze the trade-off between accuracy and efficiency. Lastly, in section 4.3, we demonstrate that our prefill and video decoding stages can be almost entirely overlapped, effectively removing a minute from the time it takes to infer an long-video input.
@@ 4.1 QuickCodec Results
Figure 2: Speed comparison of Decord, TorchCodec (with Resize), and QuickCodec on loading hour-long video. We ablate across different levels of parallelization (core counts).
@@
Video Loading Speed. We benchmark the time to load an hour-long 24 FPS 1920x1080p HD video, sampled at 1 FPS, and then resized to 448x448 pixels. The video is an hour segment of a popular movie encoded with default settings in FFmpeg using H.264, the most widely used standard [17]. Sampling frames at $1-2$ FPS is common for VideoLLMs, adopted as a trade-off between computational efficiency and understanding [4]. We resize frames to 448x448 pixels because it is the highest per frame resolution used by most VideoLLMs [48, 7]. We use an AWS m7a.16xlarge instance for our timing. Timings are averaged over five runs, all timings have a $95\%$ confidence interval of at most $0.5$ seconds. We compare QuickVideo against two video decoding frameworks designed for machine learning and integrated into existing long-video understanding inference pipelines:
@@
Decord [12]. Decord is a framework for multimedia loading designed for machine learning applications. Although it is no longer actively maintained, it remains a standard video decoding framework integrated into popular libraries like Hugging Face’s Transformers [39] and, by extension, inference libraries such as vLLM [19].
@@
TorchCodec [27]. TorchCodec is an under-development framework from the PyTorch team. It is designed to offer faster multimedia processing than TorchVision [23]. As it is a work-in-progress library, TorchCodec does not support all the features of more mature frameworks. In particular, TorchCodec does not have built-in support for frame resizing. For this reason, we report timings for the video loading and then a resizing step using TorchVision. TorchCodec is also not designed to use more than 16 cores for video decoding, we find that increasing core count beyond 16 can sometimes decrease TorchDodec’s performance.
@@
As shown in Figure 2, QuickCodec is faster than other libraries across many levels of parallelization. Whereas other frameworks plateau at 16 cores, QuickCodec scales to 64 cores. We highlight the 16 and 32 core cases, as these are the most realistic configurations. Most compute providers configure nodes with between 16 and 32 CPU cores per GPU [15, 3, 24]. Using 16-32 cores, QuickCodec is 2-3 times faster than other libraries at loading an hour-long video, reducing the time to load a video by more than 20 seconds.
@@
Speed across video durations. Our framework requires pre-computing intervals and sufficient keyframes to parallelize over. Therefore, we expect it to perform worse on shorter video durations. We benchmark QuickCodec for different video lengths, ranging from 1 minute to an hour-long video. We use the same hour-long movie video, cut to durations ranging from $1$ to $60$ minutes. We sample at 1 FPS and use $16$ cores for video decoding. All timings are averaged over 5 runs on a AWS m7a.16xlarge instance and have a $95\%$ confidence interval with at most a $0.2$ second margin of error. We find that QuickCodec is faster than other frameworks when loading videos that are more than a minute long (Figure 3). Our framework scales better in long-video loading than other libraries, increasing from 1.7x faster than Decord at loading 10 minute video to 2.1x faster at loading hour-long video. We examine additional limitations of QuickCodec compared to seek-based decoders in Appendix B.

### v2

@@ 3 Method
In this section, we introduce QuickVideo, which consists of three main components:
@@ 3.1 QuickCodec: Long Video Decoding for VideoLLMs
Given a bitstream $\mathcal{S}$ for a video $\mathcal{V}=(f_{1},f_{2},\dots,f_{m})$, where each frame has height $h$, width $w$, and the desired degree of concurrency is $c$, our goal is to compute ${\bm{F}}$ such that for all $j\in{0,1,\dots,|\mathcal{I}|-1}$, we have ${\bm{F}}_{j}=f_{\mathcal{I}[j]}$, where $\mathcal{I}\subseteq{1,2,\dots,m-1}$. In other words, ${\bm{F}}$ is a packed tensor containing all the frames selected by $\mathcal{I}$. We assume $m$ is known from container metadata or can be estimated using $pts_{max}$ and $pts_{min}$.
The efficiency of our algorithm relies on two key observations:
(1) It is faster to use $c$ cores to decode $c$ short videos than to use $c$ cores to sequentially decode a single long video. Video decoding for human playback focuses on the latter case, as humans watch earlier frames while later frames are decoded. However, due to inter-frame dependencies, sequential video decoding is difficult to parallelize [38]. In contrast, VideoLLMs require the entire video input to be loaded upfront. Therefore, we can decompose the loading of a long video $\mathcal{V}$ into loading $c$ short videos that collectively span $\mathcal{V}$. However, decoding cannot start at arbitrary frames—it must begin at keyframes. The Keyframe Intervals subroutine (Appendix A) parses the metadata of $\mathcal{S}$ and computes $c$ approximately equal-length intervals, starting and ending at keyframes, that cover the entire video. We parallelize over these intervals in Algorithm 2.
(2) VideoLLMs typically sample frames at a short, regular interval, usually 1–2 FPS [4], which is often smaller than the interval between keyframes in standard codecs. Consequently, seek-based decoding must still decode from all keyframes, leading to redundant seeks. Our algorithm requires only one seek operation per core, instead of a number of seeks proportional to the number of frames.
@@
Algorithm 2 presents the core of our video decoding algorithm. The algorithm begins by using metadata to compute $c$ keyframe-aligned intervals that span the video (line 1). Lines 2–5 initialize a shared memory block ${\bm{F}}$ and compute a dictionary $M$ that maps indices of selected frames to unique memory offsets in ${\bm{F}}$. We then decode the long video in $c$ parallel intervals (lines 6–19). Video decoding starts by seeking to the start of each interval $pts_{start}$, which is guaranteed to be a keyframe (line 7). Packets are enqueued for decoding (lines 17–18) until the decoder yields frames for processing (line 9). If the timestamp of a dequeued frame is greater than or equal to the interval endpoint $pts_{end}$, parallel processing ends (lines 11–12, 19). As the intervals in $\mathcal{J}$ span $\mathcal{S}$, $pts_{min}$ and $pts_{max}$ are given by the smallest and largest values in $\mathcal{J}$, respectively. We use Equation 1 to compute the index $i$ of $f$ (line 13). Finally, we save $f$ to ${\bm{F}}$ if $f$ is a selected frame (lines 14–16). Because decoding from a keyframe yields all frames with greater $pts$ values, and $\mathcal{D}$ outputs frames in $pts$ order, when the parallelized loop exits (line 19), all selected frames with $pts$ in the interval $[pts_{start},pts_{end})$ will have been output by $\mathcal{D}$ and saved to ${\bm{F}}$. Thus, as $\mathcal{J}$ spans $\mathcal{S}$, when the algorithm returns, ${\bm{F}}$ will contain all selected frames.
3.2 QuickPrefill: Efficient Group-Based Prefilling for VideoLLMs
@@
After decoding the video bitstream into packed tensors, they are fed into the VideoLLM for inference. However, LLM generation with long contexts is a well-known challenge due to high memory usage and computational cost. To address this, we introduce QuickPrefill, a grouped prefilling and KV cache pruning method that accelerates processing and significantly reduces memory requirements.
@@
Group-Based Prefill Let $\mathbf{X}^{v}={\mathbf{x}_{1}^{v},\ldots,\mathbf{x}_{|\mathbf{X}^{v}|}^{v}}$ denote the sequence of video tokens, where $|\mathbf{X}^{v}|$ is the total number of tokens, and each token $\mathbf{x}_{i}^{v}\in\mathbb{R}^{d}$ is a $d$-dimensional vector. To reduce memory overhead during prefilling, we partition the video token sequence into $G$ disjoint groups: $\mathbf{X}^{v}={\mathbf{X}^{v}_{1},\ldots,\mathbf{X}^{v}_{G}}$, where each group $\mathbf{X}_{g}^{v}$ contains approximately $N_{g}=\frac{|\mathbf{X}^{v}|}{G}$ tokens. Instead of processing the entire sequence at once, we sequentially prefill each group and store the corresponding key-value (KV) cache as $\mathbf{K}_{g}^{(l)},\mathbf{V}_{g}^{(l)}$ for each transformer layer $l$. This strategy significantly reduces peak activation memory usage by a factor of $G$ and remains effective even when combined with efficient attention mechanisms such as FlashAttention. Empirically, it enables hour-long video understanding while keeping GPU memory usage within practical limits (e.g., reducing memory overhead by over 100 GB; see subsection D.2).
@@
Group-Based KV Cache Pruning While group-based prefill reduces peak activation memory, the KV cache memory remains a major bottleneck. To address this, we prune unimportant KV cache vectors when processing each group, maintaining a retention ratio $\rho\in(0,1]$. This reduces KV cache memory usage by a factor of $\frac{1}{\rho}$.
The pruning decision is based on an importance score function $\mathbf{s}$, which produces an ordered list of KV entries. We select the top-$k$ KV entries until the retention ratio is satisfied:
@@
$\tilde{\mathbf{K}}_{g}^{(l)}=\mathbf{K}_{g}^{(l)}[I_{g}^{(l)}],\quad\tilde{\mathbf{V}}_{g}^{(l)}=\mathbf{V}_{g}^{(l)}[I_{g}^{(l)}],\quad\text{where }I_{g}^{(l)}=\operatorname{TopK}\left(\mathbf{s}(\mathbf{K}_{g}^{(l)},\mathbf{V}_{g}^{(l)}),;k=\rho\cdot N_{g}\right)$ (2)
@@
where $\operatorname{TopK}$ returns the indices of the top-$k$ entries. We consider several heuristic importance functions $\mathbf{s}$ from prior works [11, 16, 45]. In this paper, we primarily use the following three: 1) Key Norms (small): $\mathbf{s}=-L_{2}(\mathbf{K}_{g}^{(l)})$; 2) Value Norms: $\mathbf{s}=L_{2}(\mathbf{V}_{g}^{(l)})$; 3) Attention Scores: $\mathbf{s}=\text{matmul}(\mathbf{K}_{g}^{(l)},\mathbf{Q}^{(l)})$. Here, $L_{2}$ denotes the L2-norm function, and $\mathbf{Q}^{(l)}\in\mathbb{R}^{|\mathbf{X}^{t}|\times(n_{h}\times d_{h})}$ is the query vector of text tokens in layer $l$. We adopt Key Norms (small) as the default importance function in QuickPrefill due to its strong performance.
@@ 3.3 Overlapping QuickCodec and QuickPrefill
The preceding sections introduced two complementary components: QuickCodec for CPU-based video decoding and QuickPrefill for GPU-based group-wise prefilling. However, running these components sequentially underutilizes resources—GPUs remain idle during video decoding, and CPUs are underutilized during prefilling. To address this inefficiency, we propose an overlapped execution scheme that enables concurrent processing across CPU and GPU resources.
To achieve this, we slightly adapt frame loading: Instead of using $c$ cores to load $c$ intervals, we divide $\mathcal{V}$ into $s$ intervals, where $s\gg c$, using Keyframe Intervals($\mathcal{V},s$). We then load the frames from the $s$ intervals using $c$ cores, prioritizing earlier intervals so that frames corresponding to the first blocks of video are available sooner. This allows us to exploit QuickCodec’s fast video decoding while ensuring that early frames in $\mathcal{V}$ are prioritized for QuickPrefill. Once the frames required for the first group are loaded, QuickPrefill begins processing immediately, while QuickCodec continues decoding subsequent frames in the background. After QuickPrefill finishes processing a group, it stores the resulting KV cache and checks whether QuickCodec has loaded the frames required for the next group. If so, QuickPrefill starts processing the next group immediately. This design forms a producer-consumer pipeline between CPU decoding and GPU prefilling, ensuring the GPU is only idle if it is waiting for the CPU to finish decoding the next set of frames.
The performance improvement of this overlap strategy can be formalized. Let $t_{dec}$ and $t_{prefill}$ denote the total time for decoding and prefilling all video groups, respectively. In the sequential approach, the total execution time is $t_{dec}+t_{prefill}$. In contrast, with our overlap strategy, the execution time is:
$t_{total}=\max{t_{dec}+t^{g}_{prefill},;t_{prefill}+t^{g}_{dec}}+\Delta$ (3)
where $t^{g}_{dec}$ and $t^{g}_{prefill}$ represent the time to decode frames for the first group and to prefill the last group, respectively, and $\Delta$ is a small latency introduced by QuickCodec’s metadata parsing. Since each group contains a small number of frames and tokens, this strategy achieves near-optimal overlap between CPU and GPU resources, resulting in substantial speedup for hour-long video processing.
Note that some VideoLLMs include additional preprocessing steps (e.g., position embedding calculation or normalization), which we do not include in this analysis [4].
@@ 4 Experiments
We evaluate QuickVideo’s performance on practical long video understanding tasks. In section 4.1, we benchmark QuickCodec against existing frameworks. We also examine the limitations of QuickCodec, identifying use cases where seek-based frameworks (Algorithm 1) outperform our method. Next, in section 4.2, we evaluate the performance of QuickPrefill across four long video understanding benchmarks, analyzing the trade-off between accuracy and efficiency. Finally, in section 4.3, we demonstrate that the prefill and video decoding stages can be almost entirely overlapped, effectively reducing end-to-end inference time by nearly a minute for long video inputs.
@@ 4.1 QuickCodec Results
Figure 2: Speed comparison of Decord, TorchCodec (with Resize), and QuickCodec when loading hour-long videos. We ablate across different levels of parallelization (core counts).
@@
Video Loading Speed. We benchmark the time required to load an hour-long 24 FPS 1920x1080p HD video, sampled at 1 FPS and resized to 448x448 pixels. The video is a one-hour segment of a popular movie encoded with default FFmpeg settings using H.264, the most widely used codec [17]. Sampling frames at 1–2 FPS is a standard practice in VideoLLMs, balancing computational efficiency with task performance [4]. We resize frames to 448x448 pixels, which matches the maximum per-frame resolution used in most VideoLLMs [48, 7]. All experiments are conducted on an AWS m7a.16xlarge instance. Each timing result is averaged over five runs, with a 95% confidence interval no greater than 0.5 seconds.
We compare QuickVideo against two widely used video decoding frameworks:
@@
Decord[12]: A multimedia loading framework designed for machine learning applications. While no longer actively maintained, Decord remains integrated into popular libraries like Hugging Face’s Transformers[39] and, by extension, inference frameworks such as vLLM [19].
@@
TorchCodec[27]: A work-in-progress library from the PyTorch team designed to offer faster multimedia processing than TorchVision[23]. TorchCodec lacks some features of mature frameworks, such as built-in support for frame resizing. Thus, we report timings that combine TorchCodec loading with a resizing step via TorchVision. TorchCodec is not optimized for decoding with more than 16 cores; we observe that increasing core count beyond 16 can even degrade performance.
@@
As shown in Figure 2, QuickCodec outperforms other libraries across varying core counts. While other frameworks plateau at 16 cores, QuickCodec scales up to 64 cores. We highlight the 16- and 32-core cases as these are the most common configurations in practical deployments; most compute providers allocate between 16 and 32 CPU cores per GPU [15, 3, 24]. At 16–32 cores, QuickCodec is 2–3$\times$ faster than other libraries when loading an hour-long video, reducing video loading time by over 20 seconds.
Figure 3: Video decoding performance across different video durations (1 FPS sampling).
@@
Speed Across Video Durations. Our framework relies on pre-computing intervals and sufficient keyframes for parallelization. Therefore, we expect reduced benefits for shorter videos. We benchmark QuickCodec on videos of varying lengths, from 1 minute to 1 hour, using the same source video (an hour-long movie) cut to different durations. All tests use 1 FPS sampling and 16 cores for video decoding. Results are averaged over five runs on an AWS m7a.16xlarge instance, with a 95% confidence interval of at most 0.2 seconds. We find that QuickCodec is consistently faster than other frameworks for videos longer than 1 minute (Figure 3). Its advantage grows with video length—QuickCodec is 1.7$\times$ faster than Decord for a 10-minute video and 2.1$\times$ faster for a 1-hour video. We further discuss scenarios where seek-based decoders outperform QuickCodec in Appendix B.